### v2

# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

Abstract
The growing prevalence of generative AI inevitably raises concerns regarding the associated risks and safety implications, which catalyzes significant progress in AI safety. However, as this field thrives, a critical question emerges: Are our current efforts aligned with the broader perspective of human history and civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and create a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs. This paper identifies gaps in current approaches and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible future for AI development. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby fostering AI’s role in promoting a more secure and sustainable future for human civilization.
1 Introduction
The rapid development of AI and Large Language Models (LLMs) has fostered extensive progress in AI safety, specifically, LLM safety [171, 143, 51, 230, 90, 37]. Researchers have been dedicated to addressing potential security and privacy risks in LLM lifecycle, aiming to align AI behaviors with human values and prevent misuse of LLMs, inappropriate outputs, information leakage, etc. However, despite the intense focus on LLM safety, a critical question emerges: Are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the concerns of today?
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [146]. Despite their impressive capabilities in natural language processing and problem-solving [270, 265, 54, 38], today’s AI, including advanced LLMs [286, 260, 231], falls short of what could be considered genuine intelligence. These systems heavily rely on vast training datasets to function effectively, and lack true understanding, consciousness, self-awareness, and the ability to reason in ways comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities capable of abstract thought and genuine reasoning. Moreover, recent studies have raised concerns about the reasoning abilities of LLMs, which may actually be a form of approximate retrieval from their training data and the deductive closure of the training data [96, 239]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it fundamentally differs from human reasoning that involves abstract thinking, causal understanding, and the ability to generalize from limited examples.
The unresolved issue of sustainable energy is another critical factor that challenges our current efforts to AI safety [5]. Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [76]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent and approximately 0.02% of California’s annual electricity generation [7, 39, 169]. Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [140, 72]. Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [213], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy requirements of current AI systems raise serious sustainability concerns and limiting their deployment throughout the human society. Until we overcome the energy constraints, AI’s ability to reshape society remains restricted.
Figure 1: Illustration of Internet of Everything in a Household
These concerns suggest that our efforts on AI safety today may fail to capture deeper insights that could emerge as AI evolves, potentially neglecting technological challenges that ultimately shape the impact of AI on human civilization. This raises several critical questions:
  • Are we focusing on the right problems that align with the long-term AI development?
  • Are we potentially overlooking significant challenges that will emerge as AI evolves?
  • Are we on the right track for addressing real challenges in AI safety?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research. In this advanced future, energy issues have been resolved, and intelligent chips and brain-computer interfaces have been fully developed, which enables advanced intelligent robots to integrate into everyday life. Breakthroughs in foundational theories, such as Einstein’s relativity [57, 58] and Quantum Mechanism [59, 279], have catalyzed revolutionary progress in AI, and humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current approaches, as well as specific missions for AI safety researchers in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being constrained by these short-term solutions and look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaboration with experts from diverse fields is essential to align AI with human values better. iii) AI workflow customization and user-AI interactions are important, especially in dynamic contexts and across different user groups.
Humanity’s pursuit of higher technology will never stop. While searching for the sixpence on the ground, do not forget to look up at the night sky and look for the moon.
Paper Type Topic Notes
“Grounding and Evaluation for LLMs” [102] Survey Grounding; safety Offer valuable insights for enterprises
“Current state of risks” [8] Survey Guardrails Aspects of guardrails and challenges of implementations
“Exploring Vulnerabilities and Protections” [122] Survey Vulnerabilities Vulnerabilities of LLMs, especially prompt injections and backdoor; defenses against the attacks
“AI safety in LLMs” [37] Survey Safety issues in LLM lifecycle Safety issues towards entities in LLM lifecycle; Data/model/prompt safety; alignment; safety at scale
“Security risk” [185] Survey Attacks Attacks and risks in AI models; defensive methods
ByteDance alignment [124] Survey Broad concepts Reliability, fairness, explainability, etc
“Security of AI Agents” [82] Survey Agent security Security issues, vulnerabilities, and defenses for agents
“The good, the bad, the ugly” [272] Survey Risks; vulnerabilities; defenses LLMs as tools for security and privacy; potential risks and threats; vulnerabilities and defenses.
“Use of LLMs” [142] Survey Vulnerabilities, threats, defenses, Prevention measures, vulnerabilities, and methods to mitigate the risks associated with the misuse of LLMs
“AI risk management” [181] Insight Safety; security; governance Bridging AI safety and security in risk management; introducing a reference framework to facilitate common understanding of their differences and interplay.
“Human-ai safety” [11] Insight The integration of AI and control systems safety A unifying framework to capture safety-critical human–AI interactions; a roadmap towards next-generation human-centered AI safety
“Managing AI risks” [13] Insight Governance Emphasizing the urgent need for a balanced approach combining technical research and proactive governance to manage the extreme risks of AI systems
“Guaranteed safe AI” [43] Position Framework Define Guaranteed Safe AI to equip AI systems with formal, verifiable and auditable safety guarantees.
DeepMind ethics paper [67] Vision AI ethics Ethical challenges; interactions; impacts on society
DeepMind evaluation [252] Insight Safety evaluation Safety evaluation for genAI systems at DeepMind
Safety practices at OpenAI [151, 152, 154, 155] Blog Safety practices; framework “Lessons learned” [151], “Reimagining secure infrastructure” [152], “Preparedness Framework” [154], “OpenAI safety update” [155]
“AI governance comprehensive” [214] Book Governance AI governance framework; Case studies in industries
Table 1: Overview of AI Safety Papers
2 Related Works
Multiple AI safety surveys and benchmarks have been published in recent years. These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc. A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
Among these works, [124] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [37] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [102] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [67] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, the risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [43] proposes the concept of “Guaranteed Safe” (GS) AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior.
[11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. Specifically, the paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection. The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety research. By comparing the current landscape of AI safety with this distant future, this work identifies both the potential and limitations of today’s approaches and suggests specific missions for AI safety researchers in the 2020s. This perspective encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
3 From the New World
In this future, humanity has entered an unprecedented era. Scientists have achieved breakthroughs in energy production with technological innovations. The resulting flood of abundant, clean energy has revolutionized human civilization beyond the our ancestors’ wildest dreams, reshaping the world and initiating a new chapter of human civilization.
3.1 Imagine the Future: A Blueprint
The transformative energy revolution has catalyzed human civilization from every aspect. Foundational theories of science have undergone revolutionary breakthroughs, with major advancements in nanotechnology [136, 19, 61], Quantum Mechanics [59, 279], and Einstein’s relativity theories [57, 58]. The world has been reshaped by innovative technologies such as intelligent chips [86, 115], brain-computer interfaces [147, 81, 254], holographic technology [22, 273, 15], and advanced 3D printing [205, 73], bringing changes in daily life of miraculous.
Transportation has evolved dramatically. While simple mechanical vehicles, such as fixed-track trains and high-speed rail [144, 179, 109], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by efficient and limitless energy. Advanced autonomous vehicles and aircraft are widespread and are controlled by intelligent chips, eliminating the need for any human intervention. Also, due to breakthroughs in Einstein’s relativity theories [57, 58] and Quantum Mechanics [59, 279], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
Advanced brain-computer interfaces [147, 81, 254] and intelligent microchips [86, 115] has revolutionized the way people perceive, which expands the boundaries of human cognition. No longer are people operating devices and asking questions in search engines or LLMs; the embedded intelligence chips help them to gather more comprehensive information and interact with the surrounding environment. Meanwhile, communication over long distances has transcended the limitations of cables and the outdated internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancement in quantum computing [77, 198].
Intelligent robots have been integrated into every aspect of modern society and play important roles of daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that relies on huge amount of training data and statistical patterns in the early 21st century [131, 286], these intelligent robots perceive and learn through intelligent chips, which are real artificial intelligence and possess genuine understanding and adaptability, much like a human child discovering the world for the first time. Also, their functionalities are customizable through specialized chips based on user needs. Robots can engage in complex interactions with people and adapt their capabilities and expertise to suit their human counterparts’ requirements.
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds have been blurred, i.e., Internet of Everything. Such a network integrates objects with embedded intelligent chips, such as smart devices, autonomous systems, intelligent robots, and humans cognition, which allows for unprecedented levels of efficiency, personalization, and adaptability and fundamentally reshapes the way people interact with the world.
3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and seamlessly integrate into daily life; see Figure 1. Consequently, the concept of AI safety has been redefined to focus more on the quality of services they provide, i.e., emphasizing comprehensive quality assurance. Unlike today’s command-and-response interactions between humans and LLMs, future robots can engage in sophisticated reasoning, emotional cognition, and possess appropriate levels of autonomy. They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected system.
The Internet of Everything scenarios necessitate the robots to handle complex interactions, demanding a sophisticated understanding of the environment for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such advanced interactions require deep understanding of the environments, demanding real-time information retrieval and processing, which is far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from their environment, and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align precisely with human needs and expectations.
Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements. Even robots of the same type should be capable of adapting their services to different individuals. Teaching robots should adjust educational materials and teaching methods to capitalize on each student’s unique talents and strengths, while cleaning robots should intelligently schedule their works to minimize disruption to other humans in the household, e.g., avoiding occupied workspace. Such customization aligns the robots’ capabilities with specific user preferences, and ensures that the robots serve as effective assistants across a wide range of daily tasks and environments.
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. For instance, a care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. Conversely, while the tasks of cleaning robots are generally routine, they must still adapt to specific circumstances, such as recognizing when an area is occupied by people and postponing cleaning activities. An appropriate level of robot autonomy optimizes the division of labor between humans and robots, ensuring that robots function effectively and safely while enhancing our daily lives without compromising human control over these intelligent robots.
4 Bridging Today and the Future
Figure 2: Evolution of Human Civilization
This section presents an ambitious roadmap towards the aspirations for the advanced future, exploring key milestones in the evolution of AI and human civilization. By considering potential paradigm shifts in fundamental theories and breakthroughs in technology, we feature this fascinating journey into three eras, i.e., i) the data-driven era where AI learns knowledge from vast training datasets, ii) the neural augmentation era where human-machine interface [147, 81, 254] and intelligent chips [86, 115] extend the boundary of human cognition, and iii) the revolutionary era with a breakthrough in energy production that catalyzes transformative changes across every aspect of human society; see §4.1, §4.2, and §4.3, respectively. We then examine the evolving paradigms of AI safety throughout this journey, analyzing the potential focuses and challenges of each era and exploring the impacts of technological advancement on AI safety in §$4.$
4.1 Rookie Level: Data-Driven Era
The Data-Driven Era represents the early phase of AI development, where AI models heavily rely on the quality, availability, and quantity of training data [131]. AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [96, 239]. While AI in this era cannot be regarded as real intelligence, it assists people to solve problems and enhances productivity in the society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by diverse AI models for processing various forms of data across different scenarios [131, 50, 63, 210, 231, 260]. Researchers have been engaging in refining model structures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats. The thrive of LLMs exemplifies this progress, showcasing unprecedented capabilities of AI models in language processing and generation [260, 286, 231]. Concurrently, we observe significant advancements in other fields such as computer vision [101, 221, 256], with progress in image processing [235, 25, 216] and video analysis [223, 261, 267], as well as advancements in audio recognition [64, 178, 4, 106]. This stage is driven by exploring vast datasets and the refining algorithms continuously, which consistently pushes the boundaries of AI capabilities within specific domains.
Stage 2: Interdisciplinary AI. This stage is characterized by leveraging AI to transform traditional fields, such as healthcare [231, 2], finance [260, 282], and education [184, 135], through cross-disciplinary applications in human society. Mature multi-agent systems [116, 246, 263, 79] that leverage AI entities with different specializations will be widely utilized to assist people and solve sophisticated problems. Such agents participate in decision-making processes, or represent users to finish some tasks, providing more sophisticated solutions of a given problem. In healthcare, for example, a multi-agent system may involve collaboration between various medical agent specialties and integrate data from diverse sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will amplify the utility of AI in more complex scenarios. Although AI at this stage is still not real intelligence, it becomes deeply integrated in different domains in human society, and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
4.2 Champion Level: Neural Augmentation Era
In this era, AI models have been widely applied in human society, and humans may witness breakthroughs in advanced technologies and fundamental theories, which may extend the boundary of human cognition and life experience.
Stage 3: Breakthroughs in Advanced Technologies. In this stage, breakthroughs in advanced technologies emerge, such as brain-computer interfaces [147, 81, 254], nanotechnology [136, 19, 61], and holographic technology [22, 273, 15]. These innovations enable unprecedented levels of interaction between humans and machines by augmenting human senses and cognition. For instance, intelligent chips [86, 115] may be developed that integrate directly with the human body, creating a seamless interface between biological and digital systems. Such advances also change the nature of the data. Unlike traditional static formats such as text, images, or videos, the data in this stage becomes more comprehensive, encompassing dynamic, real-time, and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32]. In addition to the advancements in AI seen during the Data-Driven Era, this stage further expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world. However, widespread application of these techniques is limited by energy constraints and the significant resources required for implementation.
Stage 4: Breakthroughs in Fundamental Theories. In this stage, humanity experiences a new wave of breakthroughs in fundamental theories, such as Einstein’s Relativity Theories [57, 58] and Quantum Mechanics [59, 279]. These theoretical advances redefine our understanding of the universe and offer new possibilities for innovations in technologies. Such breakthroughs may also unlock solutions to longstanding challenges in human history, i.e., the energy issues, offering clean, abundant, and sustainable energy sources.
4.3 Ultimate Level: Revolutionary Era
Stage 5: Energy Revolution. A new era arrives as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation. This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [70, 200, 183], or revolutionary solar technologies [23, 137], together with advancement in nanotechnologies [136, 19, 61] that produce new materials for collecting and storing solar energy efficiently. The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space. These satellites might then transmit the abundant energy to the receivers on Earth. Regardless of the specific technology, the flood of clean, abundant energy can eliminate the dependence on fossil fuels, and fundamentally revolutionize the global economic landscapes as well as human civilization.
Stage 6: The Internet of Everything. The final stage is the realization of Internet of Everything. Every object, living being, and system becomes part of a vast, intelligent network. This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots in an area. Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures. Due to everyone is connected to a huge network, the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 273]. Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
Figure 3: Production forces of AI & core safety challenges in different historical stages.
4.4 Evolving AI Safety Paradigms
The landscape of AI safety shifts as humanity processes through stages of technology revolutions, and the focus of AI safety is different at each stage. The core AI safety challenges in different historical eras are summarized in Figure 3.
Data-Driven Era. Current AI safety practices primarily focus on data safety and model safety [37, 102, 42, 220, 167, 193, 253, 225, 224]. A critical aspect involves data privacy and security, particularly given the vast amounts of sensitive information used in AI training [118, 120, 83, 117]. AI safety researchers and practitioners also explore adversarial attacks to training data or AI models [9, 35, 119, 244, 211, 269], as well as malicious inputs designed to manipulate AI systems or induce producing unsafe content, e.g., jailbreaking or prompt injection attacks [75, 292, 126, 33, 90, 258, 293]. Correspondingly, red-teaming ML models [12, 68, 173, 197, 243] is widely used to identify potential vulnerabilities in ML models with attacks to further enhance the robustness of AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 240, 290]. Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [291] in AI systems to ensure the decision-making processes are understandable, accountable, and aligned with regulations. Other significant areas of focus is aligning AI outputs with human values [249, 174, 232, 206, 245].
Neural Augmentation Era. Advancements in cognitive chips and brain-computer interfaces have extended the boundaries of human cognition and have changed the way of human perceiving the world. With the integration of intelligent chips and their direct neural connections to AI systems, users can interact with AI with high flexibility, and the customization of AI services towards different user groups or domain knowledge might be a critical focus [248]. Individuals might rely on AI in different ways based on their background and cognitive needs, thus, personalized experiences of AI services should be explored for different user needs and preferences.
Revolutionary Era. In the revolutionary era, AI safety challenges may focus on the complex interactions between interconnected entities, such as humans and intelligent robots, within the network. As users may have diverse needs and robots will provide customized services, ensuring the safety and reliability of these interactions becomes challenging. Robots may need to request and integrate data from multiple external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Ensuring seamless communication, data integrity, and decision-making in such a sophisticated network may require smarter, more efficient methods, such as advanced cache management systems, to support the dynamic and complex interactions. This requires addressing issues of latency, data accuracy, and the security of transmitted information to maintain a safe and reliable network to ensure high-quality interactions between humans and robots in the network.
By aligning AI with the progress of human civilization, we can see the path of AI safety evolves in different stages, which reflects the growing complexity of the roles of AI in human society. Initially, the focus is on addressing safety issues related to data and models. As AI and novel technologies become more developed, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the same network.
These insights offer guidelines for what we can do for AI safety at the current stage. In the following sections, we will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed in the present stage.
5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews mainstream AI safety practices so far, and project them against the blueprint future in §3.1 to guide our current efforts in AI safety.
5.1 Jailbreaking and Red Teaming
Jailbreaking [49, 123, 175, 207, 215, 219, 234, 250, 258, 268, 293] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries utilize these techniques to manipulate the model into producing abnormal or harmful responses, such as giving instructions for illegal activities and generating offensive content. Common jailbreak methods include crafting inputs to trick the model, planting backdoors to exploiting vulnerabilities of the model, etc. Such activities pose significant risks by undermining safety measures of AI models, making it challenging to ensure the model to behaves appropriately across all scenarios.
Red-teaming is a common practice for mitigating unsafe behaviors in LLMs, where teams of experts simulate attacks or misuse scenarios to identify potential vulnerabilities in AI models [68, 243, 197, 12, 173, 128, 275]. By actively challenging AI models and probing for weaknesses through adversarial inputs, bias exploitation, or system manipulation, it helps uncover potential risks that may not be evident in normal usage, such as biases, or failure modes that might lead to undesirable behaviors in AI models. The outputs from red-teaming can be used to train the model to be less likely to cause harm or steer it away from undesirable outputs. This proactive approach has proven effective in stress-testing AI models and has been widely used in industry practices [18, 173, 68].
As AI systems grow more sophisticated, red teaming must continually adapt to keep pace with increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [153]. Also, model owners can conduct regular and diverse red teaming to anticipate and mitigate risks in evolving use cases. Moreover, current red-teaming practices are often static, typically conducted after model pretraining and consisting of one-time attacks. While this approach provides useful insights, it does not fully account for the evolving user needs, especially when dealing with customized AI services. To address this, red teaming should simulate diverse user groups and scenarios, such as simulating a diverse range of user groups and their unique needs, to address the customization of AI services and enable a comprehensive evaluation of potential vulnerabilities across different contexts.
5.2 Content Moderation and Guardrails
Content moderation involves detecting and reviewing user input and AI-generated content to ensure it complies with safety, ethical, and regulatory standards [133, 112, 80, 91]. Moderation approaches leverages rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage inputs and outputs of AI models. These approaches identify and address inappropriate content, such as toxicity, bias, hallucinations, sensitive information, and jailbreaks, to ensure that the AI models do not produce or engage with harmful content, which maintains responsible AI behaviors.
Guardrails [141, 189] employ customizable AI workflows for ensuring the quality of the model outputs while enhancing the control of the whole workflows. It functions as an intermediary layer between users and AI models, enabling users to add components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
As we reflect on practices for AI moderation and guardrails, several key challenges and considerations have emerged. The balance between utility and safety is crucial and may require dynamic solutions according to the changing contexts. Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Thus, defining appropriate thresholds for acceptable responses is challenging and requires careful consideration of the changing context and potential risks. Also, customization for moderation should be applied for different use cases and user groups, as certain information may be safe for one user group but inappropriate or dangerous for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
Topic Research Industry Practices
Jailbreaking [90, 33, 292, 258, 293, 126, 125, 170, 36, 177, 175, 219, 234, 123, 250, 49, 207, 215, 268], etc PromptArmor [227]
Red teaming [68, 243, 197, 12, 173, 128, 275, 230, 134, 48], etc Purple Llama [18], DeepMind [173], Anthropic [68], Protect AI [220], Giskard [42], Virtue AI [228], Dynamo AI [55], Mindgard [294]
Moderation [108, 259, 182, 60, 130, 92, 172, 266, 34, 107], etc OpenAI Moderation API [133], Perspective API [112], Detoxify [80], Llama Guard [91], Emergence AI [148], Giskard [42], Dynamo AI [55], Calypso AI [26], Lakera AI [226], BreezeML [24]
Guardrail - Guardrails AI [141], Nvidia Nemo Guardrails [189], Protect AI [220]
Privacy [53, 118, 139, 30, 288, 87, 281, 204], etc Dynamo AI [55], ProtectAI [220], Private AI [69], PromptArmor [227], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Zendata [229]
Defense Methods [199, 88, 240, 176, 257, 262, 133, 29, 66, 247, 278, 88, 168, 74, 40, 51, 209, 208], etc OpenAI Rule-based rewards [143]; DeepMind Weight Averaged Reward Models [188]; others see Moderation, Guardrail, and Privacy
Serving Security [258] ProtectAI [220], Relyance AI [99], Transcend [100]
AI governance [276, 277, 6, 104, 56], etc Giskard [42], ProtectAI [220], Calypso AI [26], Saidot [202], Arize [253], Dynamo AI [55], Credo AI [212], Google [187], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Transcend [100], Zendata [229], DeepMind SynthID watermark [45, 47]
Table 2: Overview of Research and Industry Practice in AI Safety Topics
5.3 AI Privacy
Privacy leakage in AI models refers to the unintended exposure of sensitive or private information that the models may have learned during training. Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [288, 87, 281]. Also, as AI models might be trained on massive datasets that could contain personally identifiable information (PII) or confidential data, they might memorize specific details of personal information, which could be inadvertently reproduced during interactions with users. Moreover, malicious users can submit crafted adversarial prompts to the AI models to extract sensitive information, e.g., jailbreaking [117, 36].
Addressing privacy leakage in AI models involves implementing methods such as differential privacy [84, 236, 285, 284], federated learning [289, 282, 78, 280, 158], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [111, 180, 95]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs. Researchers also leverage differential privacy in finetuning or in-context learning [274, 283, 222], and design privacy-preserving prompts for querying LLMs [83, 52], to maintain data confidentiality during interactions between users and AI models. Other practices for ensuring data privacy include conducting rigorous data filtering before training, real-time monitoring of model outputs (i.e., moderation in §5.2), etc, such that the model outputs would not reveal sensitive information [133, 112, 80, 91] .
As we reflect on the practices of data privacy, a crucial aspect is the balance between privacy and utility. While ensuring data confidentiality is essential, it often comes at the cost of reduced functionality and utility. In practice, implementing privacy measures in AI services may not effectively prevent malicious users but inadvertently harm the experiences for benign users. This highlights a deeper issue: those determined to exploit vulnerabilities of AI services often find new, sophisticated ways to bypass safeguards, whereas being overly cautious may impact regular users negatively.
5.4 AI Security and Defense Methods
Defense methods typically involve the following strategies to ensure that AI models are robust and aligned with ethical guidelines: (1) employing prompt engineering methods, together with chain-of-thought reasoning or multi-agent systems, to help models better understand user inputs and reject harmful requests or unsafe outputs [176, 257, 247, 278]; (2) using built-in methods like exploring the decoding stage [88, 51], leveraging reward-based approaches [143, 188], or utilizing hidden states [29, 51] to prevent the generation of unsafe content; (3) constructing datasets with safety features, such as secure instructions or adversarial samples and training models with this data to improve their robustness against harmful inputs [262, 66, 133, 10, 168]; (4) aligning models with human values through Supervised Fine-Tuning (SFT) [287] and Reinforcement Learning from Human Feedback (RLHF) [157], which helps models to follow user instructions safely and adhere to desired ethical norms; (5) training models to detect unsafe content and developing guardrail pipelines to prevent models from taking unsafe inputs or generating inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [74, 40]; and (7) utilizing differential privacy  [1, 93, 209], homomorphic encryption [218, 95], secure MPC [44, 105], to ensure that sensitive data is protected during model training, etc.
5.5 AI Serving Security
AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI models when serving users [162, 258]. It focuses on protecting models and interactions from various threats, such as such as supply chain vulnerabilities [164], model theft [161], model denial of service [160], and insecure plugins [159].
Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers, or user prompts collected with the supply chain [164]. Such vulnerabilities include using outdated models, deprecated third-party packages, or improper handling of training data, leading to data breaches, malicious injections, or even system failures [164]. Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [165]. Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe outputs of LLMs [163].
Model theft involves unauthorized access, replication, or reverse engineering of LLM weights and parameters to create functionally equivalent copies [161], posing risks to intellectual property, brand reputation, and financial security. Distributed denial of service (DDoS) attacks can compromise AI services, especially when models are integrated into widely accessible applications [160]. Attackers can craft prompts that exploit recursive behaviors in LLMs, which consumes excessive computational resources [190]. For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, which affected ChatGPT and developer tools for hours in November 2023 [166].
Plugins enable interactions with external software, databases, web tools, or APIs, which enhances LLM functionality [159]. However, they may introduce vulnerabilities if proper execution controls are lacking. Adversaries can exploit insecure plugins with adversarial prompts to execute unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [159]. Without proper input validation, plugins may reveal sensitive information [191]. Moreover, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst case, might cause the underlying system to crash [191, 192].
5.6 AI Governance
AI governance is a system of rules, practices, processes, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values, fulfilling legal standards, and meeting ethical principles [132, 21]. [214] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) Establish Accountability for AI; 2) Assess Regulatory Risks; 3) Gather Inventory of Use Cases; 4) Increase Value of Underlying Data; 5) Assess Fairness and Accessibility; 6) Improve Reliability and Safety; 7) Heighten Transparency and Explainability; interpretability of AI. 8) Implement Accountability with Human-in-the-Loop; 9) Support Privacy and Retention; 10) Improve Security; 11) Implement AI Model Lifecycle and Registry; 12) Manage Risk; 13) Realize AI Value.
Among these directions, 1) and 8) are related to accountability; 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (§5.1), content moderation and guardrails (§5.2), AI privacy (§5.3), AI security and defense methods (§5.4), and AI serving security (§5.5); 3) is related to business and is out of the scope of this subsection (see [214] for details on practical AI governance use cases); 4) is related to data management such as training data processing [133], access control [203, 156, 16, 6], and data regulations [241, 14]; 5) is related to customizing AI workflows, i.e., guardrails (§5.2) and access control in AI systems [97, 6]; 7) is related to AI transparency; and 11) and 13) are related to customizing AI workflow for enhanced functionality and safety monitoring to the AI systems, i.e., guardrail (§5.2). Without loss of generality, this subsection discusses the rest topics, including i) AI accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
Accountability. Accountability refers to being responsible for actions and impacts of AI systems on individuals and society. It ensures responsibility and auditability, enables redress, and minimizes negative impact of AI systems [138]. However, defining and achieving accountability is challenging due to AI’s complexity and various considerations like privacy and intellectual property [187, 138]. [149] defines accountability as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power. It outlines a framework with seven features, including context, range, agent, forum, standards, process, and implications, as well as four key goals, including compliance, reporting, oversight, and enforcement. These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance. Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve the goal of accountability [149]. Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers, and across various sectors such as government, private industry, and academia [186]. Approaches include algorithmic assessments, auditing, and data provenance techniques [186]. Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [276, 277].
Regulations. AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI technologies. European Union’s AI Act [238] categorizes AI applications by risk level, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk. High-risk AI applications, like those in healthcare and law enforcement, must meet high requirements of transparency, fairness, and safety. The General Data Protection Regulation (GDPR) [237] enforces strict data protection and privacy requirements. The Blueprint for an AI Bill of Rights [85] and NIST AI Risk Management Framework [150] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
Access control. Access control in AI systems manages user access to specific information within the model according to their permissions, thereby limiting unauthorized data exposure and ensuring compliance with data privacy regulations. Effective methods include data access controls [97] and model access controls [6]. As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for.
Interpretability, explainability, and transparency. Interpretability, explainability, and transparency are interrelated concepts but differ in their scope and focus. Interpretability focuses on understanding how an AI model functions internally and produces its outputs [124, 71, 194, 181]. Explainability provides explanations to specific AI outputs, presenting the results in ways that can be easily understood and trusted by end-users [37, 124, 71, 181, 264]. Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how decisions are made, what data is used, and why specific results are produced [255, 110, 71, 181]. Transparency also extends to properties like traceability and content provenance that tracks models and data to ensure accountability and reliability [214, 45, 47]. In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [291, 56, 71, 181, 233]. Practices to improve transparency, interpretability, and explainability include a variety of methods, including i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [41, 129, 124, 129, 201, 217, 46], counterfactual explanations [242, 98], concept activation explanations [103, 124], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [251, zhang2022automatic,diao2023active, 65], tree-of-thought [127, 271, 28], graph-of-thought [17, 113], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [276, 277, 45, 47]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [114, 94], function calling [62, 121], or web browsing [145], as reference documents for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate the interpretability of AI systems [195, 196, 20].
6 AI Safety in the 2020s: Challenges and Missions
This section utilizes the blueprint to guide our efforts today, navigating the challenges and opportunities of the current stage. As AI safety practitioners in the 2020s, we stand at a pivotal moment in history, with society moving toward an era where AI will be deeply integrated into daily life. While addressing current safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape.
A shift from AI safety to AI quality assurance. The rapid changing AI landscape requires a shift of focus from narrowly defined safety concerns to a more comprehensive concept, i.e., AI quality assurance. Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [157] of model outputs as well as the quality of interactions between humans and AI to enhance the overall quality of AI services. Viewing safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications. The terminology “quality assurance” also guide us to think more about what we can do to align our current efforts with the blueprint picture.
Alignment with more interdisciplinary collaborations. A key area that demands increased focus is the alignment of AI with deeper insights across different fields, people, and cultures. Current AI models, while advanced, often lack a deep understanding of the ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field. However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from different fields such as psychology, sociology, history, art, and anthropology, etc, especially for AI applications for specific fields such as finance [260, 282], healthcare [231, 2], education [184, 135], etc. AI safety practitioners must work towards bridging this gap by fostering interdisciplinary collaborations. By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as the ethical, cultural, and social dynamics of our society.
Hallucination. Hallucination remains a fundamental and persistent challenge in AI systems, and is difficult to eliminate entirely, even as AI advances [89]. This is because hallucinations can be traced back to nearly every stage in the lifecycle of AI models, such as the quality and representativeness of training data, inappropriate user inputs, and the processes used during inference. Low-quality or biased training data can cause models to generate incorrect or misleading information, while inappropriate user inputs may lead to the creation of false details. During inference, models can also produce outputs that deviate from factual accuracy when they attempt to fill gaps in their training knowledge. Despite implementing Retrieval-Augmented Generation (RAG) [114] can mitigate hallucination, it remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage. Achieving this requires an ongoing effort to refine training data, ground user queries properly, and implement moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
Customization. Customization in AI workflow is essential for ensuring the AI systems to meet the diverse user needs while operating effectively in a variety of contexts. Current AI systems such as ChatGPT [133] tend to provide same services across all user groups, regardless of individual differences. However, as AI becomes more integrated into everyday life, one-size-fits-all solutions are no longer adequate. People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and relevant services. As an example, an AI model interacting with children should respond differently compared with interacting with adults, which requires to take into account age, learning objectives, cognitive abilities, emotional maturity, etc. Customization of AI services can enhance safety by adjusting contexts and interactions to align with specific cultural norms, legal regulations, and ethical considerations. In fields such as finance [260, 282], healthcare [231, 2], and education [184, 135], customizing AI services based on user profiles is crucial for preventing undesirable outputs and enhancing user experiences.
Interaction. From our long-term goal towards the blueprint future, the advanced society involves complex interactions between intelligent robots and human users (similar ideas are also mentioned in [11, 67]). This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment at the current stage. Besides prompt-based LLM systems or services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to maintain coherent and multi-round interactions that adapts to evolving conversational contexts. Correspondingly, new safety protocols are in need to adjust in real-time to evolving contexts and user needs. It is also essential to develop new methodologies for safety assessment and mitigation that can function effectively in complex, rapidly evolving environments with multiple AI agents and human users.
7 Conclusion
This paper presents a blueprint for an advanced human society and leverages this forward-looking vision to guide today’s AI safety efforts. Through the blueprint, it becomes clear that Artificial General Intelligence (AGI) is not the ultimate goal of AI development. Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into every aspect of life. While AGI is a popular topic discussed among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields. Regarding AGI as the ultimate goal might limit our creativity in the 2020s, trapping us in “local optima” and potentially causing us to overlook the real challenges.
What will this future world be like? Will it bring more happiness to human-beings? The answer is uncertain. Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability in everyday life. On the other hand, with many everyday problems being solved by advanced technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
As we stand at a pivotal moment in human history, our efforts become not only a technological endeavor but also a profound exploration of human potential and the boundaries of human capability. In any case, the journey toward this advanced future is fascinating, and every human being will look forward to it.